# Constraint Back-translation Improves Complex Instruction Following of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) struggle to follow instructions with complex constraints in format, length, etc. Following the conventional instruction-tuning practice, previous works conduct post-training on complex instruction-response pairs generated by feeding complex instructions to advanced LLMs. However, even advanced LLMs cannot follow complex instructions well, thus limiting the quality of generated data. In this work, we find that existing datasets inherently contain implicit complex constraints and propose a novel data generation technique, constraint back-translation. Specifically, we take the high-quality instruction-response pairs in existing datasets and only adopt advanced LLMs to add complex constraints already met by the responses to the instructions, which naturally reduces costs and data noise. In the experiments, we adopt Llama3-70B-Instruct to back-translate constraints and create a high-quality complex instruction-response dataset, named Crab. We present that post-training on Crab improves multiple backbone LLMs’ complex instruction-following ability, evaluated on extensive instruction-following benchmarks. We further find that constraint back-translation also serves as a useful auxiliary training objective in post-training. Our code, data, and models will be released to facilitate future research. [^1] https://github.com/THU-KEG/Crab
@@ 2.1 Constructing Crab
We then adopt the Llama3-70B-Instruct LLM and Python scripts to back-translate constraints for the seed data. First, we define a set of $19$ types of common constraints, such as length and style constraints. Considering some constraints, e.g., length constraint, cannot be effectively followed by LLMs (Sun et al., 2024), leading to noisy back-translation, and some constraints can be easily generated using Python scripts, we choose to adopt Python scripts for $6$ types of constraints, including length, keyword, and punctuation constraints. Specifically, we write and paraphrase several templates for each of these $6$ types of constraints. For example, for a length constraint, one template is “Please generate a response with fewer than <placeholder> words but more than <placeholder> words”. We then use Python scripts to automatically identify the value for this constraint in the response and fill the templates to construct a constraint. For the length constraint, we randomly sample a range that includes the value to fill the template. For keyword and punctuation constraints, we randomly select corresponding items present in the response to fill the templates of constraints. For the remaining $13$ types of constraints, such as style constraints, we use Llama3-70B-Instruct for constraint back-translation. Specifically, we include these $13$ types of constraints and corresponding descriptions, along with the instruction and response, as the prompt to query Llama3-70B-Instruct to generate the constraints implicitly satisfied by the response. We then use Llama3-70B-Instruct to re-verify whether the response satisfies the generated constraints and exclude the constraints that are not met. We adopt ROUGE-L (Lin, 2004) with a threshold of $0.6$ to exclude constraints that are similar to the instruction and other constraints. Finally, we randomly sample and manually review $50$ instruction-response pairs and their generated constraints, finding minimal noise and high compliance between constraints and response.
@@ 3.1 Experimental Setup
We adopt two widely-used open-source base models, Mistral 7B (Jiang et al., 2023) and Llama 3 8B (Dubey et al., 2024), as our backbone models for developing Llama3Crab and MistralCrab. Specifically, we employ Mistral-7B-v0.3 and Meta-Llama-3-8B, downloaded from Hugging Face (Wolf et al., 2019). During the SFT stage, we adopt a $5\times 10^{-6}$ learning rate, $256$ batch size, and train the Mistral for $4$ epochs and Llama 3 for $3$ epochs. During the DPO optimization stage, we adopt $5\times 10^{-7}$ learning rate, $64$ batch size, and $1$ training epoch.
@@
Our baselines include popular open-source and proprietary LLMs, divided into three main categories for comparison: (1) Proprietary LLMs, including GPT-3.5 (OpenAI, 2022) and GPT-4 (OpenAI, 2024). (2) General instruction-tuning LLMs, including Vicuna-V1.5 13B (Chiang et al., 2023), trained on the 125k ShareGPT dataset, WizardLM-V1.2 13B (Xu et al., 2023), trained on the 196k Evol-Instruct dataset, and Zephyr beta 7B (Tunstall et al., 2023), trained with the UltraFeedback (Cui et al., 2023) dataset using the DPO objective (Rafailov et al., 2023), which achieves leading performance on chat benchmarks based on Mistral 7B. (3) Models specifically optimized for complex instruction-following tasks. We use the Conifer series (Sun et al., 2024), which first generates constraints based on seed instructions and then generates corresponding responses based on instructions and constraints using GPT-4, and achieves the state-of-the-art performance in complex instruction-following at the 7B model scale.
@@ 3.5 Analysis on Constraint Category
Figure 5: Experimental results on different categories of constraints in FollowBench of MistralCrab and ConiferSFT.
@@ 4.1 Instruction Following
Instruction following involves following user intentions to generate helpful responses, which is fundamental to modern LLMs (Zhang et al., 2023). Ouyang et al. (2022) first propose the practice of aligning LLMs to follow human instructions, using SFT and RLHF to train models, which is the key factor in the success of ChatGPT (OpenAI, 2022). Subsequently, numerous studies focus on enhancing the instruction-following capabilities of LLMs, particularly for open-source models, which can be summarized in two main aspects: (1) data-driven approaches, which design an automated pipeline or use human annotation to produce high-quality training data (Xu et al., 2023; Taori et al., 2023; Chiang et al., 2023; Peng et al., 2023; Kim et al., 2023; Cui et al., 2023; Mukherjee et al., 2023; Ivison et al., 2023; Köpf et al., 2024; Qi et al., 2024; Liu et al., 2024; Li et al., 2024; Bai et al., 2024a). (2) new training methods, including novel objectives (Rafailov et al., 2023; Gallego, 2024; Zhou et al., 2024; Hejna and Sadigh, 2024; Meng et al., 2024) or training pipelines (Tunstall et al., 2023; Li et al., 2024; Yuan et al., 2024; Chen et al., 2024).
A more challenging instruction following scenario is complex constrained instruction following, where the responses should further satisfy specific constraints, such as length. Previous studies have shown that LLMs struggle to follow these instructions (Jiang et al., 2024; Qin et al., 2024). Recent efforts focus on enhancing this ability by constructing high-quality training data (Sun et al., 2024; He et al., 2024). This process typically involves collecting a set of seed instructions, constructing constraints, and then generating responses based on the instructions and constraints using advanced LLMs. In this work, we introduce constraint back-translation, which directly generates constraints from existing instruction and response pairs, which reduces the generation costs and noise.
@@ 4.2 Back-translation
Recently, several studies have explored applying back-translation to the field of large language models to efficiently generate high-quality data automatically (Li et al., 2023a; Pham et al., 2024; Köksal et al., 2023). Li et al. (2023a) proposed reversing the training objective to automatically generate corresponding instructions for existing unsupervised corpora, while Pham et al. (2024) and Köksal et al. (2023) leveraged the powerful general capabilities of LLMs to generate instructions from the corpus directly. These works usually focus on instruction back-translation, which generate instructions from existing text corpora. In this work, we propose constraint back-translation, which generates high-quality constraints based on instructions and responses. It can effectively generate constraints without altering the original responses, and reduces data noise and the cost of data generation.
@@ 5 Conclusion
We discuss potential ethical concerns related to this work: (1) Intellectual property. Our research leverages several widely used SFT datasets, and we strictly comply with their licensing terms. We will share Crab under the CC BY-SA 4.0 license [^2] https://creativecommons.org/licenses/by-sa/4.0/. (2) Intended use and Potential risk control. The goal of this paper is to introduce Crab, designed to enhance the performance of LLMs on complex instruction tasks. Crab is built using widely available public datasets. We trust that the original publishers have anonymized and sanitized these datasets appropriately. Additionally, we randomly sampled $100$ instances and found no sensitive information. (3) AI assistance. We used GPT-4 to paraphrase some sentences and check grammar.
@@ Appendix A Data Collection
In this section, we provide a detailed explanation of our data construction process, divided into two parts: the details of constraint construction (§ A.1) and the data distribution of Crab (§ A.2).
@@ A.1 Details of Constraints Construction
Table 4 presents all $19$ types of constraints we defined. It is important to note that for the "Define circumstances", clarifying the subject or object, or defining the circumstances under which the instruction applies, we observed that generating this constraint independently often results in this additional constraint being too similar to the original instruction. Therefore, we integrate it directly with the original interaction to develop a refined instruction. If selected during the combination process, instead of being added to the instruction like other constraints, it replaces the original instruction. Since not all constraints are applicable to every output text (e.g., Special Output Format refers to the special format of the output, such as Python, tables, JSON, HTML, or LaTeX), we apply weighted sampling during the combination process to ensure a more balanced sampling of all constraints.
Among the constraints calculated using Python scripts, two categories are particularly unique: (1) Number-related categories: such as Length and Words Per Sentence, where we used NLTK (Loper and Bird, 2002) for calculation. (2) Keyword: We applied the lightweight, unsupervised keyword extraction method Yake (Campos et al., 2020) to extract the top $3$ most significant keywords from the output text.
A.2 Dataset Distrubution
@@ Appendix B Model Training
Constraint Category Generator Weighted
Define Circumstances LLM 0.7
Desired Writing Style LLM 0.7
Semantic Elements LLM 0.8
Morphological Constraints LLM 0.8
Multi-lingual Constraints LLM 0.8
Specific Literary Devices LLM 0.8
Specific Grammatical Structure LLM 0.8
Hierarchical Instructions LLM 0.8
Special Output Format LLM 0.8
Paragraphs Constraints LLM 0.7
Specific Sentence LLM 0.7
Keyword Formatting LLM 1.0
Item Listing Details LLM 1.0
Length Constraint Python 0.5
Words Per Sentence Python 0.5
Sentences Per Paragraph Python 0.3
Characters Per Word Python 0.3
Keyword Constraint Python 0.5
Punctuation Limitation Python 0.3
Table 4: Constraint types defined during the back-translation process. "Weighted" indicates the sampling probability of each constraint type during combination.

### v2

@@ Abstract
Large language models (LLMs) struggle to follow instructions with complex constraints in format, length, etc. Following the conventional instruction-tuning practice, previous works conduct post-training on complex instruction-response pairs generated by feeding complex instructions to advanced LLMs. However, even advanced LLMs cannot follow complex instructions well, thus limiting the quality of generated data. In this work, we find that existing datasets inherently contain implicit complex constraints and propose a novel data generation technique, constraint back-translation. Specifically, we take the high-quality instruction-response pairs in existing datasets and only adopt advanced LLMs to add complex constraints already met by the responses to the instructions, which naturally reduces costs and data noise. In the experiments, we adopt Llama3-70B-Instruct to back-translate constraints and create a high-quality complex instruction-response dataset, named Crab. We present that post-training on Crab improves multiple backbone LLMs’ complex instruction-following ability, evaluated on extensive instruction-following benchmarks. We further find that constraint back-translation also serves as a useful auxiliary training objective in post-training. Our code, data, and models will be released to facilitate future research [^1] https://github.com/THU-KEG/Crab.
@@ 2.1 Constructing Crab
We adopt the Llama3-70B-Instruct LLM and Python scripts to back-translate constraints for the seed data. We ultimately use Llama3-70B-Instruct to automatically generate constraints implicitly satisfied by the response from instruction-response pairs. To enhance the diversity of generated constraints, we manually collect $13$ commonly used constraint types [^2] We sample $200$ instruction-response pairs from ShareGPT to observe real-world constraint needs and summarize them. as examples in the prompt for constraint generation, which results in over 100 constraint types. We then use Llama3-70B-Instruct to re-verify whether the response satisfies the generated constraints and exclude the constraints that are not met. Considering some constraints, e.g., length constraint, cannot be effectively followed by LLMs (Sun et al., 2024), leading to noisy back-translation, and some constraints can be easily generated using Python scripts, we choose to adopt Python scripts for $6$ types of constraints. Specifically, we write and paraphrase several templates for each of these constraints. For example, for a length constraint, one template is “Please generate a response with fewer than <placeholder> words but more than <placeholder> words”. We then use Python scripts to automatically identify the value for this constraint in the response and fill the templates to construct a constraint. For the length constraint, we randomly sample a range that includes the value to fill the template. For keyword and punctuation constraints, we randomly select corresponding items present in the response to fill the templates of constraints. We adopt ROUGE-L (Lin, 2004) with a threshold of $0.6$ to exclude similar constraints. Finally, we sample and manually review $50$ instruction-response pairs and their generated constraints, finding minimal noise and high compliance between constraints and response.
@@ 3.1 Experimental Setup
We adopt two widely-used open-source base models, Mistral 7B (Jiang et al., 2023) and Llama 3 8B (Dubey et al., 2024), as our backbone models for developing Llama3Crab and MistralCrab. Specifically, we employ Mistral-7B-v0.3 and Meta-Llama-3-8B, downloaded from Hugging Face (Wolf et al., 2019). During the SFT stage, we adopt a $5\times 10^{-6}$ learning rate, $256$ batch size, and train the Mistral for $4$ epochs and Llama 3 for $3$ epochs. During the DPO optimization stage, we adopt $5\times 10^{-7}$ learning rate, $64$ batch size, and $1$ training epoch with beta $0.01$ for Mistral and $3$ epochs with beta $0.1$ for Llama 3.
@@
Our baselines include popular open-source and proprietary LLMs, divided into three main categories for comparison: (1) Proprietary LLMs, including GPT-3.5 (OpenAI, 2022) and GPT-4 (OpenAI, 2024). (2) General instruction-tuning LLMs, including Vicuna-V1.5 13B (Chiang et al., 2023), trained on the 125k ShareGPT dataset, WizardLM-V1.2 13B (Xu et al., 2023), trained on the 196k Evol-Instruct dataset, Zephyr beta 7B (Tunstall et al., 2023), trained with the UltraFeedback (Cui et al., 2023) dataset using the DPO objective (Rafailov et al., 2023), and Mistral-Instruct 7B v0.3 (Jiang et al., 2023), which achieves leading performance on chat benchmarks based on Mistral 7B (3) Models specifically optimized for complex instruction-following tasks, including Suri-I-ORPO (Pham et al., 2024), which is optimized for multi-constraint instruction-following tasks in long-form text generation, and the Conifer series (Sun et al., 2024), which are trained on the data where the constraints, instructions, and responses are all generated using GPT-4.
@@ 3.5 Analysis on Constraint Category
Figure 5: Experimental results on different categories of constraints in the FollowBench of MistralCrab and ConiferSFT.
@@ 4.1 Instruction Following
Instruction following task involves following user intentions to generate helpful responses, which is fundamental to modern LLMs (Zhang et al., 2023). Ouyang et al. (2022) first propose the practice of aligning LLMs to follow human instructions, using SFT and RLHF to train models, which is the key factor in the success of ChatGPT (OpenAI, 2022). Subsequently, numerous studies focus on enhancing the instruction-following capabilities of LLMs, particularly for open-source models, which can be summarized in two main aspects: (1) data-driven approaches, which design an automated pipeline or use human annotation to produce high-quality training data (Xu et al., 2023; Taori et al., 2023; Peng et al., 2023; Chiang et al., 2023; Mukherjee et al., 2023; Qi et al., 2024; Liu et al., 2024; Li et al., 2024; Bai et al., 2024a; Peng et al., 2024). (2) new training methods, including novel objectives (Rafailov et al., 2023; Gallego, 2024; Zhou et al., 2024; Hejna and Sadigh, 2024; Meng et al., 2024) or training pipelines (Tunstall et al., 2023; Li et al., 2024; Yuan et al., 2024; Chen et al., 2024b).
A more challenging instruction following scenario is complex constrained instruction following, where the responses should further satisfy specific constraints, such as length, keyword, and format. Previous studies have shown that LLMs struggle to follow these instructions (Jiang et al., 2024; Qin et al., 2024; Chen et al., 2024a). Recent efforts focus on enhancing this ability by constructing high-quality training data (Sun et al., 2024; He et al., 2024). This process typically involves collecting a set of instructions, constructing constraints, and then generating responses based on the instructions and constraints using advanced LLMs. This work introduces constraint back-translation, which generates constraints from instruction-response pairs, reducing data construction costs and noise.
@@ 4.2 Back-translation
Recently, several studies have explored applying back-translation to the field of large language models to efficiently generate high-quality data automatically (Li et al., 2023a; Pham et al., 2024; Köksal et al., 2023). Li et al. (2023a) proposed reversing the training objective to automatically generate corresponding instructions for existing unsupervised corpora, while Pham et al. (2024) and Köksal et al. (2023) leveraged the powerful general capabilities of LLMs to generate instructions from the corpus directly. Although Pham et al. (2024) also generated constraints, it fell within the realm of instruction back-translation and did not involve dedicated optimization or exploration for constraint generation. In this work, we propose constraint back-translation, an effective data generation approach that generates high-quality constraints based on instruction-response pairs.
@@ 5 Conclusion
We discuss potential ethical concerns related to this work: (1) Intellectual property. Our research leverages several widely used SFT datasets, and we strictly comply with their licensing terms. We will share Crab under the CC BY-SA 4.0 license [^3] https://creativecommons.org/licenses/by-sa/4.0/. (2) Intended use and Potential risk control. The goal of this paper is to introduce Crab, designed to enhance the performance of LLMs on complex instruction tasks. Crab is built using widely available public datasets. We trust that the original publishers have anonymized and sanitized these datasets appropriately. The data construction process does not include additional social bias. Additionally, we randomly sampled $100$ instances from our dataset and found no sensitive information. (3) AI assistance. We used GPT-4 to paraphrase some sentences and check grammar.
@@ Appendix A Data Collection
In this section, we provide a detailed explanation of our data construction process, divided into three parts: the details of constraint construction (§ A.1), the data diversity of Crab (§ A.2) and the data distribution of Crab (§ A.3).
@@ A.1 Details of Constraints Construction
Table 4 presents the definitions of each constraint in our constraint set. It is important to note that for the “Situation”, clarifying the subject or object, or defining the circumstances under which the instruction applies, we observed that generating this constraint independently often results in this additional constraint being too similar to the original instruction. Therefore, we integrate it directly with the original interaction to develop a refined instruction. If selected during the combination process, instead of being added to the instruction like other constraints, it replaces the original instruction.
Among the constraints calculated using Python scripts, two categories are particularly unique: (1) Number-related categories: such as Length and Words Per Sentence, where we used NLTK (Loper and Bird, 2002) for calculation. (2) Keyword: We applied the lightweight, unsupervised keyword extraction method Yake (Campos et al., 2020) to extract the top $3$ most significant keywords from the output text. Table 9 provides an example generated after the constraint back-translation process.
A.2 Dataset Diversity
We adopted $4$ widely used post-training datasets for constructing our CRAB dataset: Alpaca GPT-4, Open Assistant, Evol-Instruct, and Orca Chat. These datasets are of high quality and contain diverse data instances for generating rich constraints. The “Rate” column in Table 4 shows the distribution of constraint types in the CRAB dataset. Moreover, for constraints generated by the LLM, since we provided $13$ types of constraints as examples in the prompt, we analyzed the keywords of each generated constraint and extracted the top $10$ keywords for each major category, presenting them as subcategories in Figure 7. The “situation” category was not included in the analysis because it is task-specific, making clustering difficult.
A.3 Dataset Distrubution
@@ Appendix B Model Training
Constraint Category Description Generator Rate
Situation Adding conditions, clarifying the subject or object, or defining the circumstances under which the instruction applies. LLM 36.9
Writing Style Specify the style requirements for the response to align with the intended message and audience. LLM 81.3
Semantic Elements Clearly articulate the main theme, focus, meaning, or underlying concept of the response. LLM 99.5
Morphological Outline specific prohibitions, such as avoiding certain words or phrases and refraining from specific formatting styles. LLM 99.7
Multi-lingual Specify the language(s). LLM 94.8
Literary Devices Identify any particular literary devices to be employed. LLM 91.7
Grammatical Structure Specify the grammatical structure. LLM 99.1
Hierarchical Instructions Establish a response hierarchy, defining the prioritization and structuring of tasks within the output. LLM 83.1
Output Format Depending on the required format of the output—such as Python, tables, JSON, HTML, LaTeX—impose relevant format constraints. LLM 15.2
Paragraphs Constraints Clearly specify the required number of paragraphs or sections in the text. Additionally, indicate any specific spacing or separators needed—such as blank lines, horizontal rules, or special symbols to enhance readability and visual appeal. LLM 71.3
Specific Sentence Specify a particular phrase to be included either at the beginning or end of the text, clearly indicating its exact placement. LLM 70.1
Header Format Specify the formatting style for titles or keywords within the Output, such as using bold, italics, or CAPITAL LETTERS. LLM 9.5
Item Listing Details Clearly specify the formatting for individual entries within the text. Direct the use of specific symbols for listing—such as bullet points (•), numbers (1., 2., 3., etc.), or hyphens (-). LLM 67.7
Length Constraint Determine the word count of the output text to establish length constraints. Python 47.6
Word Constraint Determine the number of words in each sentence to set word constraints. Python 15.1
Sentence Constraint Determine the number of sentences in each paragraph to establish sentence constraints. Python 20.6
Character Constraint Determine the number of characters in each word. Python 21.4
Keyword Constraint Determine the keywords in the output text to make the constraints more detailed. Python 45.1
Punctuation Limitation Specify which punctuation marks cannot be used in the output text. Python 21.6
Table 4: Constraint types defined during the back-translation process. The Rate (%) indicates the proportion of instances in the entire dataset that generated constraints of this category.
@@ Appendix C Details on the impact of constraints on output quality
Appendix D More Results
In this section, we present additional experimental results, divided into five parts: full results on Followbench (§ D.1), a fairer comparison where Conifer is replaced with the same backbone as ours (§ D.2), our experimental results on LLaMA3.2-3B (§ D.3), the performance of general tasks (§ D.4), a discussion about the effect of the effect of syntactic constraints (§ D.5) and a discussion about reverse training (§ D.6).
D.1 SSR results on Followbench
We report FollowBench results under the Hard Satisfaction Rate (HSR) metric in Table 1. Table 5 presents results on FollowBench under Soft Satisfaction Rate (SSR) metric. We also conducted a comparison with the ShareGPT version, which is trained exclusively on the ShareGPT dataset.
Model Backbone IFEval FollowBench (SSR) AVG
[S]P [S]I [L]P [L]I AVG L1 L2 L3 L4 L5 AVG
Llama3-ShareGPT* Llama3 $23.7$ $26.4$ $33.8$ $37.1$ $30.3$ $44.0$ $40.0$ $39.6$ $33.3$ $33.6$ $38.1$ $34.2$
Llama3Crab Llama3 $39.4$ $50.2$ $43.8$ $54.2$ $46.9$ $57.5$ $52.4$ $51.2$ $47.0$ $45.6$ $50.7$ $48.8$
Llama3Crab + DPO Llama3 $40.3$ $52.0$ $47.7$ $58.9$ $49.7$ $64.6$ $55.8$ $54.7$ $52.4$ $54.0$ $56.3$ $53.0$
[…]
MistralCrab Mistral $47.9$ $57.3$ $51.6$ $61.2$ $54.5$ $63.9$ $60.6$ $55.1$ $50.4$ $49.4$ $55.9$ $55.2$
MistralCrab + DPO Mistral $49.7$ $61.5$ $57.7$ $68.5$ $59.4$ $66.1$ $59.2$ $59.8$ $55.3$ $51.2$ $58.3$ $58.8$
Table 5: Full results (%) on IFEval and FollowBench, where $\dagger$ and * indicate that the results are sourced from  Sun et al. (2024) and  Dong et al. (2024), respectively.
Model Backbone IFEval FollowBench (HSR) AVG
[S]P [S]I [L]P [L]I AVG L1 L2 L3 L4 L5 AVG
[…]
ConiferSFT-7B-v0.3 Mistral $45.8$ $57.0$ $49.7$ $60.8$ $53.3$ $60.6$ $52.2$ $46.7$ $38.8$ $26.5$ $45.0$ $49.1$
[…]
ConiferDPO-7B-v0.3 Mistral $46.4$ $57.2$ $54.9$ $64.6$ $55.8$ $60.1$ $52.5$ $46.6$ $45.7$ $38.6$ $48.7$ $52.2$
MistralCrab + DPO Mistral $49.7$ $61.5$ $57.7$ $68.5$ $59.4$ $66.1$ $59.2$ $59.8$ $55.3$ $51.2$ $58.3$ $58.8$
Llama3.2 3B Llama3.2 $15.0$ $26.3$ $15.5$ $26.7$ $20.9$ $12.7$ $14.7$ $14.9$ $18.2$ $11.8$ $14.5$ $17.7$
Llama3.2Crab Llama3.2 $34.9$ $44.6$ $38.1$ $48.1$ $41.4$ $51.7$ $36.4$ $29.1$ $19.7$ $14.3$ $30.2$ $35.8$
Table 6: Experimental results (%) of the original Conifer paper, our reproduced results on Mistral 7B v0.3 (the backbone used in MistralCrab) and the results of Llama3.2 as the backbone for IFEval and FollowBench. Here, $\dagger$ indicates that the results are from Sun et al. (2024).
Model MMLU GSM8k MathQA ARC challenge ARC easy Hellaswag SIQA AVG
Mistral-7B-v0.3 $62.3$ $37.2$ $35.4$ $48.8$ $79.5$ $60.9$ $46.0$ $52.9$
MistralCrab $59.8$ $46.6$ $34.4$ $47.4$ $77.5$ $62.5$ $47.7$ $53.7$
MistralCrab + DPO $61.0$ $34.0$ $35.7$ $53.2$ $79.8$ $68.1$ $51.0$ $54.6$
Table 7: Experimental results (%) on general benchmarks.
Model IFEval FollowBench (HSR) AVG
[S]P [S]I [L]P [L]I AVG L1 L2 L3 L4 L5 AVG
MistralCrab (w/o syntactic constraints) $44.2$ $54.2$ $47.5$ $57.9$ $50.9$ $55.6$ $50.1$ $33.9$ $25.1$ $17.8$ $36.5$ $43.7$
MistralCrab (mixing throughout the process) $46.0$ $56.4$ $49.9$ $60.1$ $53.1$ $60.6$ $49.1$ $36.5$ $29.4$ $15.6$ $38.2$ $45.7$
MistralCrab (keeping them separate) $47.9$ $57.3$ $51.6$ $61.2$ $54.5$ $63.9$ $54.4$ $40.1$ $30.4$ $27.9$ $43.3$ $48.9$
Table 8: Experimental results (%) of different mixing strategies and the effect of syntactic constraints.
D.2 Different Backbone Comparison
Since  Sun et al. (2024) did not specify the model version of Conifer, we reproduced Conifer on Mistral-7B-v0.3, which is the backbone used in MistralCrab, and the results are presented in Table 6. All conclusions remain consistent with those stated in the main text.
D.3 More Backbone Models
To verify the scalability and applicability of our approach, we conducted experiments using LLaMA3.2-3B (Dubey et al., 2024) as the backbone, with all training hyperparameters consistent with those of LLaMA3-8B. As shown in Table 6, CRAB significantly improves the model’s performance on complex instruction-following tasks.
D.4 Results on General Benchmarks
We evaluated our model on a general benchmark. We select several widely-used benchmarks for assessing general capabilities: MMLU (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), MathQA (Amini et al., 2019) and Commonsense Reasoning (including HellaSwag (Zellers et al., 2019), SIQA (Sap et al., 2019), ARC easy and challenge (Clark et al., 2018)). As shown in Table 7, our models improve the average performance of the original Mistral model. Especially on commonsense reasoning tasks (ARC challenge, ARC easy, HellaSwag, SIQA), MistralCrab + DPO achieved an average improvement of $4.25\%$, showing that enhanced instruction-following boosts reasoning performance.
D.5 Discussion about the effect of syntactic constraints
Syntactic constraints, such as dependency relations, are common in benchmarks and user queries. Including them in our dataset improves the model’s instruction-following ability, as evidenced by the performance drop when these constraints are removed, as shown in Table 8.
D.6 Discussion about Reverse Training
Intuitively, reverse training enhances the model’s understanding of constraints by requiring it to generate them, as this process inherently demands a deeper comprehension. We chose to keep forward and reverse data separately following Berglund et al. (2023), which separates data in their “both” setting. We also experimented with mixing both types of training throughout the process, and the results shown in Table 8 indicate that keeping them separate leads to better model performance. A possible reason may be that mixing the forward and reverse training may lead to overfitting easily.
Figure 7: The Top-$10$ subcategories for $12$ major constraint types (excluding "situation") generated by the LLM after constraint back-translation, as the "situation" type is closely tied to the input and lacks clustering tendencies.